given a 10% fall in petroleum price and a 6.2 percent rise in daily wage paid to hired labour used for the production of rice in ghana. estimate the net effect of these changes on the production of rice, given that the elasticities production due petroleum price and daily wage are 1.15 and 1.25. respectively.


1
Expert's answer
2022-04-12T13:22:33-0400

0,1*1,15 = 0,115

0,062*1,25 = 0,0775

0,115 - 0,0775 = 0,0375

So it means, that price of the product will decrease on 3,75%